# FeatSharp: Your Vision Model Features, Sharper

### v1

@@ Abstract
The feature maps of vision encoders are fundamental to myriad modern AI tasks, ranging from core perception algorithms (e.g. semantic segmentation, object detection, depth perception, etc.) to modern multimodal understanding in vision-language models (VLMs). Currently, in computer vision, the frontier of general purpose vision backbones are Vision Transformers (ViT), typically trained using contrastive loss (e.g. CLIP). A key problem with most off-the-shelf ViTs, particularly CLIP, is that these models are inflexibly low resolution. Most run at 224x224px, while the “high resolution” versions are around 378-448px, but still inflexible. We introduce a novel method to coherently and cheaply upsample the feature maps of low-res vision encoders while picking up on fine-grained details that would otherwise be lost due to resolution. We demonstrate the effectiveness of this approach on core perception tasks as well as within agglomerative model (RADIO) training as a way of providing richer targets for distillation.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Upsampling architecture diagram. We combine the upsampled features coming from FeatUp (Fu et al., 2024) with the tiled features and mix them with FeatSharp to produce a feature map with higher fidelity. The tiled features have more detail, but also representation issues such as the difference in upper and lower body at the tile boundary. “Full-Res Reference” is for display purposes, as for a model that doesn’t exhibit stable resolution scaling (e.g. DFN CLIP, SigLIP, etc.) we don’t have access to a target hi-res feature map. Our proposed learnable modules are in green, combined with a learnable FeatUp JBU stack.
@@
The use of vision foundation models (VFM) (Awais et al., 2023) has seen widespread use since the beginning of the modern era of computer vision using deep learning (Krizhevsky et al., 2012), primarily used to perform transfer learning (Plested & Gedeon, 2022) (e.g. finetuning a VFM on a downstream task), information retrieval (Babenko et al., 2014; Zhang & Liu, 2024), and most recently, to power visual capabilities for vision-language models (VLM) (Alayrac et al., 2022; et al., 2024; Liu et al., 2023; Lin et al., 2023). The recent shift toward using Transformers (Vaswani et al., 2017) for computer vision (ViT (Dosovitskiy et al., 2021)) has both tremendously moved the field forward, but has generally left the use of VFMs in a tricky spot: Transformers are computationally demanding and have poor algorithmic scaling properties ($O(n^{2})$ for 1D sequences, or $O(\left(w\cdot h\right)^{2}$ for 2D inputs), leaving the majority of models to be relatively low-resolution. For example, perhaps the most popular family of VFMs to date, CLIP (Radford et al., 2021), typically runs at 224 or 336px input resolutions, and produces spatial features at a 14x downsample (e.g. $224^{2}\rightarrow 16^{2}$). Owing to the nature of learned position embeddings, ViTs also tend to be relatively inflexible to changes of input resolution, allowing for changes, but requiring finetuning (Dosovitskiy et al., 2021).
@@
It’s possible that the strict dependence on the training resolution is an artifact of the algorithm used for training, as DINOv2 (Oquab et al., 2023; Darcet et al., 2023) is quite robust to interpolating its position embeddings, producing stable features at various resolutions (Ranzinger et al., 2024b), ignoring for the moment that DINOv2, being a transformer, is expensive to use at high-resolution. A recent technique called AM-RADIO (Ranzinger et al., 2024a), borrowing ideas from ViTDet (Li et al., 2022), FlexiViT (Beyer et al., 2023), and RO-ViT (Kim et al., 2023), has attempted to create a resolution-flexible ViT, however it is still dependent on low-resolution ViTs as it distills from other seminal VFMs which are low-res-only: DFN CLIP (Fang et al., 2023) and SigLIP (Zhai et al., 2023).
@@
Recently, FeatUp (Fu et al., 2024) aims to directly address the problem of low-resolution vision features by using one of two learned upsampling algorithms: A model-specific generalized upsampler using Joint Bilateral Upsampling (JBU) (Kopf et al., 2007), or a model-specific-image-specific implicit network. While they demonstrate particularly compelling results with their implicit network, their results using the stack of JBU filters lack refined details (shown in figure 18 in the appendix). Along with lack of granular refinement, it’s impossible for this approach to capture fine-grained details that are too small for the vision backbone to detect at its native resolution. To this end, we take inspiration from both FeatUp’s JBU approach, as well as the recent trend in VLMs such as LLaVA 1.6 (Liu et al., 2024), InternVL-1.5 (Chen et al., 2024), NVLM (Dai et al., 2024b) and Eagle (Shi et al., 2024b) to tile an image, aggregating local features from a fixed-low-resolution model, to build an upsampler that simultaneously leverages the raw pixel guidance, low-res feature guidance, and regional tile guidance, resulting in substantially more detailed feature maps which are also capable of capturing details too small for the original resolution. Specifically, we:
@@
  • Build on top of FeatUp’s JBU algorithm (Fu et al., 2024) by adding de-biasing and tiling fusion modules to incorporate detailed tile features, resulting in significantly higher levels of detail, with extensive experiments demonstrating effectiveness
@@
  • Study the relationship between FeatUp’s feature consistency and ViT-Denoiser’s (Yang et al., 2024a) approach to cleaning the features of a ViT at its native resolution
@@
  • Introduce an improved training setting for AM-RADIO (Ranzinger et al., 2024a) demonstrating a +$0.6\%$ improvement across the entire benchmark suite, and better teacher adapter features
@@ 2 Related Work
Figure 2: Diagram of the FeatSharp module. We first concatenate the bilinear upsampled and tiled mosaic feature maps along the channel dimension. We then apply a transformer block with sliding window attention followed by MLP (in this case, SwiGLU), and then slice off the first half of the channels, corresponding to the bilinear upsampled buffer. The role of FeatSharp thus is to refine the bilinear buffer by leveraging the tile buffer.
@@
The most obvious baseline for feature upsampling is to use traditional filtering approaches such as bilinear or bicubic upsampling. The alternative is to evaluate the network at higher resolution, however it comes with the dual drawback that computational cost increases (quadratically in the case of Vision Transformers), and also that many models (ViTs in particular) have trouble extrapolating from their trained resolution (Beyer et al., 2023; Dehghani et al., 2023). If we expand our view to include parametric approaches, then deconvolution (Noh et al., 2015; Shi et al., 2016; Dumoulin & Visin, 2016) and resize-conv (Odena et al., 2016) are popular choices. There are also pixel-adaptive approaches such as CARAFE (Wang et al., 2019), SAPA (Lu et al., 2022), and FeatUp (Fu et al., 2024).
@@
The use of tiling to increase information is currently very prominent in VLMs (Liu et al., 2024; Chen et al., 2024; Dai et al., 2024a), albeit an alternative approach is to instead leverage the models at hi-res themselves (Beyer et al., 2024; Wang et al., 2024). We also see RADIO-Amp(Heinrich et al., 2024) being primarily useful at high-resolution within VLMs. In the increasingly VLM-centric approach to computer vision, we turn our focus to RADIO-Amp, as it has a training procedure that relies on matching a high-resolution student against a low-resolution teacher, an application area that is perfect for studying feature upsampling, as it would provide richer guidance to the distillation.
@@ 3.2 Feature Normalization
FeatUp supports either leaving the features coming from the backbone as-is (e.g. no normalization), or using a LayerNorm to better condition the outputs for feature learning. For a similar motivation as PHI-S (Ranzinger et al., 2024a), we want to avoid using the raw features as they have varying spreads, and we’d also like to avoid using LayerNorm as it makes the features incompatible with the original feature space. Naïvely learning the raw feature space across the suite of teachers without normalization often led to convergence issues, particularly given the wide variance of activations.
Tile-Guided Attentional Refinement
@@
Joint-Bilateral Upsampling is able to retain object boundaries primarily in instances when there are noticeable changes in intensity in the RGB input image. This results in sharp contours, but within a region, we end up with vague and blurry feature representations. Owing to the reliance on raw pixel intensities, object contours that are less discriminative in color space often get blurred with the neighborhood. Finally, because the upsampling operation is only truly operating on the low resolution feature maps of the model, it’s impossible for JBU to introduce new details into the feature map that are visible/encodable at higher input resolutions. FeatUp’s implicit upsampler doesn’t have this problem because it’s constructing a unified view from numerous local views of the original image, enabling detailed refinements. We propose an intermediary method between JBU which leverages a single featurizer inference, and the implicit model, which relies on numerous inferences and is thus cost prohibitive [^1] https://github.com/mhamilton723/FeatUp/issues/2.
@@
Inspired by the use of tiling in Vision-Language Models (VLMs) (Liu et al., 2024; Shi et al., 2024a; Dai et al., 2024b), we develop an attentional refinement block that is able to integrate the information between a bilinear upsampled feature map, as well as a feature map composed of tiles. We show an overview of the algorithm in figures 1, 2 and 3. The diagram shows actual results using RADIOv2.5-L, which is the most scale equivariant foundation model (Heinrich et al., 2024), and generally the strongest visual foundation model (Lu et al., 2024; Drozdova et al., 2024; Guo et al., 2024). Because the model has strong resolution scaling, it provides us with a good way to compare the results of the upsampling process against the feature maps of the same resolution attained by increasing the resolution of the input image. We also observe that even just at 4×4$\times$4 × tiling, there are major discontinuities in the tiled feature map, which the FeatSharp module must overcome to produce a unified higher resolution image.
@@
For the FeatSharp module, we leverage a single Attention+SwiGLU transformer block. In order to prevent the quadratic cost of global attention, we instead use local attention (Ramachandran et al., 2019). We concatenate the JBU upsampled buffer with the tiled feature map and feed it to the block. After the block is computed, we slice the first $C$ dimensions of the output, with $C$ being the model feature dimension, and treat that as the refined features. The slicing strategy takes advantage of the fact that a transformer block has a residual pathway, and thus a no-op from the transformer would be equivalent to returning the bilinear upsampled features. Through the attention mechanism, the model is able to consider the local neighborhood and refine its features to achieve better multi-view consistency. To this end, we train our model identically to FeatUp’s multi-view consistency algorithm. We do not employ any special loss functions beyond the MSE loss on multi-view consistency, contrary to FeatUp’s use of Total Variation and Conditional Random Field losses. We provide ablations wrt architecture choice in appendix B.2.
@@ 3.4 Complexity
An important point about this method is that because of the tiling, it requires more evaluations of the base vision model to construct the high resolution feature map. However, due to the scaling properties of global self-attention, our proposed method always has better scaling properties than running the original model at higher resolution (assuming the model is capable of doing this in the first place):
$\begin{split}f(x)&=c\left(1+x^{2}\right)\\ g(x)&=c\left(x^{2}\right)^{2}=cx^{4}\\ f(x)&\leq g(x)\quad\forall x>1\end{split}$ (6)
where $f(x)$ is the relative cost of computing FeatSharp with $x\in\mathbb{Z}_{+}$ upsample factor, $g(x)$ is the cost of the hi-res image based on quadratic scaling on the number of patches (and thus tiles), and $c$ being the cost of processing a single tile. We show the actual scaling cost in figure 15 in the appendix.
@@ 4.2 Qualitative
We run this upsampling method on seven different foundation models coming from diverse domains such as supervised (ViT, (Dosovitskiy et al., 2021)), contrastive (DFN CLIP (Fang et al., 2023), SigLIP (Zhai et al., 2023)), Self-supervised (DINOv2-L-reg (Darcet et al., 2023)), Segmentation (SAM (Kirillov et al., 2023)), VLM (PaliGemma (Beyer et al., 2024)), and Agglomerative (RADIOv2.5-L (Ranzinger et al., 2024a)). Results are in figure 6. The original feature maps run the spectrum from extremely noisy (SigLIP) to very clean (RADIOv2.5-L, SAM), which allows us to demonstrate the effectiveness of the approach on a diverse set of models. Taking SAM for an example, the way in which is has thick edge outlines cannot be reproduced in the shape interior by FeatUp, primarily because the bilateral upsampler is operating on the raw pixels, where the interior edge doesn’t exist in the real image. For all of the featurizers, FeatSharp is able to achieve more legible representations, particularly it’s more able to closely match the real hi-res features in the second column.
Figure 6: PCA visualizations of features from a basketball scene. Column 1: Raw features produced by the model at normal resolution (e.g. 14x downsample for DFN CLIP, SigLIP, PaliGemma, and DINOv2, 16x downsample for SAM and RADIOv2.5-L. Column 2: Raw features at the 4x upsample resolution (we interpolate the position embeddings for those models that don’t natively support resolution changes). Column 3: FeatUp 4x upsampling (prior work). Column 4: FeatSharp 4x upsampling.

### v2

@@ Abstract
The feature maps of vision encoders are fundamental to myriad modern AI tasks, ranging from core perception algorithms (e.g. semantic segmentation, object detection, depth perception, etc.) to modern multimodal understanding in vision-language models (VLMs). Currently, in computer vision, the frontier of general purpose vision backbones is Vision Transformers (ViT), typically trained using contrastive loss (e.g. CLIP). A key problem with most off-the-shelf ViTs, particularly CLIP, is that these models are inflexibly low resolution. Most run at $224\times 224$px, while the “high-resolution” versions are around $378-448$px, but still inflexible. We introduce a novel method to coherently and cheaply upsample the feature maps of low-resolution vision encoders while picking up on fine-grained details that would otherwise be lost due to resolution. We demonstrate the effectiveness of this approach on core perception tasks as well as within agglomerative model training using RADIO as a way of providing richer targets for distillation. Code available at https://github.com/NVlabs/FeatSharp.
Machine Learning, ICML, Computer Vision, Feature Upsampling
@@ 1 Introduction
Figure 1: PCA visualizations of features from a basketball scene. Column 1: Raw features produced by the model at normal resolution (e.g. 14x downsample for DFN CLIP, SigLIP, PaliGemma, and DINOv2, 16x downsample for SAM and RADIOv2.5-L. Column 2: Raw features at the 4x upsample resolution (we interpolate the position embeddings for those models that don’t natively support resolution changes). Column 3: FeatUp-JBU 4x upsampling (prior work). Column 4: FeatSharp 4x upsampling.
@@
Figure 2: Upsampling architecture diagram. We combine the upsampled features coming from FeatUp (Fu et al., 2024) with the tiled features and mix them with FeatSharp to produce a feature map with higher fidelity. The tiled features have more detail, but also representation issues such as the difference in upper and lower body at the tile boundary. “Full-Res Reference” is for display purposes, as for a model that doesn’t exhibit stable resolution scaling (e.g. DFN CLIP, SigLIP, etc.) we don’t have access to a target hi-res feature map. Learned modules have a fire icon, and frozen modules a snowflake.
@@
Vision foundation models (VFM) (Awais et al., 2023) have seen widespread use since the beginning of the modern era of computer vision using deep learning (Krizhevsky et al., 2012), primarily used to perform transfer learning (Plested & Gedeon, 2022) (e.g. finetuning a VFM on a downstream task), information retrieval (Babenko et al., 2014; Zhang & Liu, 2024), and most recently, to power visual capabilities for vision-language models (VLM) (Alayrac et al., 2022; et al., 2024; Liu et al., 2023; Lin et al., 2023). The recent shift toward using Transformers (Vaswani et al., 2017) for computer vision (ViT (Dosovitskiy et al., 2021)) has tremendously moved the field forward, but has generally left the use of VFMs in a tricky spot: Transformers are computationally demanding and have poor algorithmic scaling properties ($O(n^{2})$ for 1D sequences, or $O(\left(w\cdot h\right)^{2}$ for 2D inputs), leaving the majority of models to be relatively low-resolution. For example, perhaps the most popular family of VFMs to date, CLIP (Radford et al., 2021), typically runs at 224 or 336px input resolutions, and produces spatial features at a 14x downsample (e.g. $224^{2}\rightarrow 16^{2}$). Owing to the nature of learned position embeddings, ViTs also tend to be relatively inflexible to changes of input resolution, allowing for changes, but requiring finetuning (Dosovitskiy et al., 2021).
@@
It is possible that the strict dependence on the training resolution is an artifact of the algorithm used for training, as DINOv2 (Oquab et al., 2023; Darcet et al., 2023) is quite robust to interpolating its position embeddings, producing stable features at various resolutions (Ranzinger et al., 2024b), ignoring for the moment that DINOv2, being a transformer, is expensive to use at high-resolution. A recent technique called AM-RADIO (Ranzinger et al., 2024a), borrowing ideas from ViTDet (Li et al., 2022), FlexiViT (Beyer et al., 2023), and RO-ViT (Kim et al., 2023), has attempted to create a resolution-flexible ViT, however it is still dependent on low-resolution ViTs as it distills from other seminal VFMs which are low-res only: DFN CLIP (Fang et al., 2023) and SigLIP (Zhai et al., 2023).
@@
Recently, FeatUp (Fu et al., 2024) aims to directly address the problem of low-resolution vision features by using one of two learned upsampling algorithms: A model-specific generalized upsampler using Joint Bilateral Upsampling (JBU) (Kopf et al., 2007), or a model-specific-image-specific implicit network. While they demonstrate particularly compelling results with their implicit network, their results using the stack of JBU filters lack refined details (shown in figure 17 in the appendix). Along with the lack of granular refinement, it’s impossible for this approach to capture fine-grained details that are too small for the vision backbone to detect at its native resolution. To this end, we take inspiration from both FeatUp’s JBU approach, as well as the recent trend in VLMs such as LLaVA 1.6 (Liu et al., 2024), InternVL-1.5 (Chen et al., 2024), NVLM (Dai et al., 2024b) and Eagle (Shi et al., 2024b) to tile an image, aggregating local features from a fixed-low-resolution model, to build an upsampler that simultaneously leverages the raw pixel guidance, low-res feature guidance, and regional tile guidance, resulting in substantially more detailed feature maps which are also capable of capturing details too small for the original resolution. Specifically, we:
@@
  • Build on top of FeatUp’s JBU algorithm (Fu et al., 2024) by adding de-biasing and tiling fusion modules to incorporate detailed tile features, resulting in significantly higher levels of detail, with extensive experiments demonstrating effectiveness.
@@
  • Study the relationship between FeatUp’s feature consistency and ViT-Denoiser’s (Yang et al., 2024a) approach to cleaning the features of a ViT at its native resolution.
@@
  • Introduce an improved training setting for AM-RADIO (Ranzinger et al., 2024a) demonstrating a +$0.6\%$ improvement across the entire benchmark suite, and better teacher adapter features.
@@ 2 Related Work
Figure 3: Diagram of the FeatSharp module. We first concatenate the JBU upsampled and tiled mosaic feature maps along the channel dimension. We then apply a transformer block with sliding window attention followed by MLP (in this case, SwiGLU), and then slice off the first half of the channels, corresponding to the bilinear upsampled buffer. The role of FeatSharp thus is to refine the JBU buffer by leveraging the tile buffer.
@@
The most obvious baseline for feature upsampling is to use traditional filtering approaches such as bilinear or bicubic upsampling. The alternative is to evaluate the network at higher resolution, however it comes with the dual drawback that computational cost increases (quadratically in the case of Vision Transformers), and also that many models (ViTs in particular) have trouble extrapolating from their trained resolution (Beyer et al., 2023; Dehghani et al., 2023). If we expand our view to include parametric approaches, then deconvolution (Noh et al., 2015; Shi et al., 2016; Dumoulin & Visin, 2016) and resize-conv (Odena et al., 2016) are popular choices. There are also pixel-adaptive approaches such as CARAFE (Wang et al., 2019), SAPA (Lu et al., 2022), LiFT (Suri et al., 2024), and FeatUp (Fu et al., 2024).
@@
The use of tiling to increase information is currently very prominent in VLMs (Liu et al., 2024; Chen et al., 2024; Dai et al., 2024a), albeit an alternative approach is to instead leverage the models at hi-res themselves (Beyer et al., 2024; Wang et al., 2024). We also see RADIOv2.5(Heinrich et al., 2024) being primarily useful at high-resolution within VLMs. In the increasingly VLM-centric approach to computer vision, we turn our focus to RADIOv2.5, as it has a training procedure that relies on matching a high-resolution student against a low-resolution teacher, an application area that is perfect for studying feature upsampling, as it would provide richer guidance to the distillation.
@@ 3.2 Feature Normalization
FeatUp supports either leaving the features coming from the backbone as-is (e.g. no normalization), or using a LayerNorm to better condition the outputs for feature learning. For a similar motivation as PHI-S (Ranzinger et al., 2024a), we want to avoid using the raw features as they have very distinct distributions, and we’d also like to avoid using LayerNorm as it makes the features incompatible with the original feature space. Naïvely learning the raw feature space across the suite of teachers without normalization often led to convergence issues, particularly given the wide variance of activations. Thus, we adopt PHI-S as a way to standardize the backbone features without distortion and to retain the ability to model the original distribution. We compute the distribution statistics over 100k samples from the training dataset.
3.3 Tile-Guided Attentional Refinement
@@
Joint-Bilateral Upsampling is able to retain object boundaries primarily in instances when there are noticeable changes in intensity in the RGB input image. This results in sharp contours, but within a region, we end up with vague and blurry feature representations. Owing to the reliance on raw pixel intensities, object contours that are less discriminative in color space often get blurred with the neighborhood. Finally, because the upsampling operation is only truly operating on the low resolution feature maps of the model, it’s impossible for JBU to introduce new details into the feature map that are visible/encodable at higher input resolutions. FeatUp’s implicit upsampler doesn’t have this same limitation because it’s constructing a global view from numerous local views of the original image, enabling detailed refinements. We propose an intermediary method between JBU which leverages a single featurizer inference, and the implicit model, which relies on numerous inferences and is thus cost prohibitive [^1] https://github.com/mhamilton723/FeatUp/issues/2.
@@
Inspired by the use of tiling in Vision-Language Models (VLMs) (Liu et al., 2024; Shi et al., 2024a; Dai et al., 2024b), we develop an attentional refinement block that is able to integrate the information between a JBU upsampled feature map, as well as a feature map composed of tiles. We show an overview of the algorithm in figures 2, 3 and 4. The diagram shows actual results using RADIOv2.5-L, which is the most scale equivariant foundation model (Heinrich et al., 2024), and generally the strongest visual foundation model (Lu et al., 2024; Drozdova et al., 2024; Guo et al., 2024). Because the model has strong resolution scaling, it provides us with a good way to compare the results of the upsampling process against the feature maps of the same resolution attained by increasing the resolution of the input image. We also observe that even just at 4×4$\times$4 × tiling, there are major discontinuities in the tiled feature map, which the FeatSharp module must overcome to produce a unified higher-resolution image.
@@
For the FeatSharp module, we leverage a single Attention+SwiGLU transformer block. In order to prevent the quadratic cost of global attention, we instead use 2D local attention (Ramachandran et al., 2019). We concatenate the JBU upsampled buffer with the tiled feature map and feed it to the block. After the block is computed, we slice the first $C$ dimensions of the output, with $C$ being the model feature dimension, and treat that as the refined features. The slicing strategy takes advantage of the fact that a transformer block has a residual pathway, and thus a no-op from the transformer would be equivalent to returning the bilinear upsampled features. Through the attention mechanism, the model is able to consider the local neighborhood and refine its features to achieve better multi-view consistency. To this end, we train our model identically to FeatUp’s multi-view consistency algorithm. We do not employ any special loss functions beyond the MSE loss on multi-view consistency, contrary to FeatUp’s use of Total Variation and Conditional Random Field losses. We provide ablations wrt architecture choice in appendix B.2.
@@ 3.5 Complexity
An important point about this method is that because of the tiling, it requires more evaluations of the base vision model to construct the high-resolution feature map. However, due to the scaling properties of global self-attention, our proposed method always has better scaling properties than running the original model at higher resolution (assuming the model is capable of doing this in the first place). Specifically, let $f(x)$ be the relative cost of computing FeatSharp, and $g(x)$ the relative cost of running the base model on the hi-res input, with $x\in\mathbb{Z}_{+}$ being the number of tiles per dimension, and $c$ being the cost of processing a single tile:
$\begin{split}f(x)&\leq c\sum_{i=1}^{x}i^{2}\\ g(x)&=c\left(x^{2}\right)^{2}=cx^{4}\\ f(x)&\leq g(x)\quad\forall x>1\end{split}$ (6)
We show the empirical scaling cost in figure 14 in the appendix, and prove equation 6 in appendix E.2. We also note that experiments for FeatSharp only use the global view, plus the final level of tiles, thus $f(x)$ simplifies to $f(x)=c\left(1+x^{2}\right)$, however we prove the general case, as progressive upsampling may be beneficial in future work.
@@ 4.2 Qualitative
We run this upsampling method on seven different foundation models coming from diverse domains such as supervised (ViT, (Dosovitskiy et al., 2021)), contrastive (DFN CLIP (Fang et al., 2023), SigLIP (Zhai et al., 2023)), Self-supervised (DINOv2-L-reg (Darcet et al., 2023)), Segmentation (SAM (Kirillov et al., 2023)), VLM (PaliGemma (Beyer et al., 2024)), and Agglomerative (RADIOv2.5-L (Ranzinger et al., 2024a)). Results are in figure 1. The original feature maps run the spectrum from extremely noisy (SigLIP) to very clean (RADIOv2.5-L, SAM), which allows us to demonstrate the effectiveness of the approach on a diverse set of models. Taking SAM for an example, the way in which it has thick edge outlines cannot be reproduced in the shape interior by FeatUp, primarily because the bilateral upsampler is operating on the raw pixels, where the interior edge doesn’t exist in the real image. For all of the featurizers, FeatSharp is able to achieve more legible representations. In particular it is more able to closely match the real hi-res features in the second column.